OpenStudy (anonymous):

MEDAL!!!HELP!!! Using the following equation, find the center and radius of the circle. You must show all work and calculations to receive credit. x2 + 2x + y2 + 4y = 19 (8 points)

OpenStudy (anonymous):

Help???

ganeshie8 (ganeshie8):

start by grouping x and y terms separately

OpenStudy (anonymous):

Oops the first one is x^2 And y2 is actually y^2

ganeshie8 (ganeshie8):

\[\large x^2 + 2x + y^2 + 4y = 19\] \[\large (x^2 + 2x) + (y^2 + 4y) = 19\]

ganeshie8 (ganeshie8):

like that, eh ?

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

So, now what do I do?

ganeshie8 (ganeshie8):

heard of the phrase `completing the square` before ?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Wait... yeah I did

ganeshie8 (ganeshie8):

good, next step is to complete the square for the stuff inside each parenthesis

OpenStudy (anonymous):

So I add one to both sides

OpenStudy (anonymous):

Then I factor the parentheses???

ganeshie8 (ganeshie8):

kindof...

OpenStudy (anonymous):

Are you still on?

ganeshie8 (ganeshie8):

\[\large x^2 + 2x + y^2 + 4y = 19\] \[\large (x^2 + 2x) + (y^2 + 4y) = 19\] \[\large (x^2 + 2x + \color{red}{1^2}) + (y^2 + 4y +\color{red}{ 2^2}) = 19 + \color{Red}{1^2 + 2^2}\]

OpenStudy (anonymous):

Hello?

ganeshie8 (ganeshie8):

we're almost done, knw how to complete the square ?

ganeshie8 (ganeshie8):

use below identity : \(a^2 + 2ab + b^2 = (a+b)^2\)

OpenStudy (anonymous):

I heard it but I forgot

OpenStudy (anonymous):

What;s that mean?

ganeshie8 (ganeshie8):

\[\large (x^2 + 2x + \color{red}{1^2}) + (y^2 + 4y +\color{red}{ 2^2}) = 19 + \color{Red}{1^2 + 2^2} \] \[\large (x+1)^2 + (y+2)^2 = 19 + 1 + 4 \] \[\large (x+1)^2 + (y+2)^2 = 24 \]

ganeshie8 (ganeshie8):

compare that equation with the standard form of circle : \[\large (x-h)^2 + (y-k)^2 = r^2\] center = \(\large (h, k)\) radius = \(\large r\)